If a cycle wheel of radius 4 m completes one revolution in two seconds, then acceleration of a point on the cycle wheel will be-(1)π^2m/s^2(2)2π^2m/s^2(3)4π^2m/s^2(4)8π^2m/s^2

Grade:11

1 Answers

Vikas TU
7 years ago
angular velocity would be:
w = theta/time = 2*pi/2 = pi rad
accln. = w^R = > pi^2 * 4 = 4*pi^2 m/s^2.